VOLTAGE CONTROLLED GEOMETRIC PHASE ROTATION IN ^{208}Pb^{19}F.

\underline{J. E. FURNEAUX}, NEIL SHAFER-RAY, J. COKER, P. M. RUPASINGHE,, Homer L. Dodge Department of Physics and Astronomy, University of Oklahoma, Norman, OK 73019; and C. P. McRAVEN, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973.

Many theoretical publications have investigated the impact of the geometric phase on measurements of the e-EDM. However, there has been surprisingly little quantitative comparison of these models with experiment. Here we create a quantum beat experiment that starts with an optical pump and ends with an optical probe of^{208}Pb^{19}F. This measurement includes the ability to control a geometric phase variation of the molecular alignment by applying an appropriate bias voltage. These experiments will then used to test the accuracy of our model calculations of geometric phase rotation.